# AdTEC: A Unified Benchmark for Evaluating Text Quality in Search Engine Advertising

### v1

@@ 5 Result and Discussion
Table 6 provides an overview of the result. For the complete result including base size encoder models and GPT-3.5, please refer to Appendix Table 10.
@@
Fine-tuned Setting with Encoder Models. XLM-RoBERTa and Tohoku BERT achieved the highest or competitive scores in two or more tasks. In addition, the LARGE model outperformed the BASE model in most tasks, suggesting that the increasing parameter size plays a key role in understanding ad expressions.
@@
Zero-/Few-shot Setting with LLMs. GPT-4 achieved high performance across all three tasks, while ELYZA7b performed the best among the LLMs in Ad Consistency. A substantial difference was observed between the open and OpenAI’s LLMs in the Ad Similarity task. CALM27b and ELYZA7b scored close to 0, indicating no correlation, whereas OpenAI’s models, especially GPT-4, achieved competitive scores of 0.776/0.811. Thus, the open LLMs used in this study struggle with handling the semantic similarity or numerical answers, whereas GPT-4 performs considerably better. In the Ad Performance Estimation task, all LLMs produced uncorrelated responses close to 0, indicating that accurately predicting ad performance remains a challenge for open LLMs.
@@
Fine-tuned Encoder Models vs. Zero-/Few-shot LLMs. Overall, the fine-tuned models outperformed the LLMs. The difference was substantial, ranging from 0.2 to 0.6, especially for the Ad Acceptability, Ad Performance Estimation, and A3 Recognition tasks. This can be attributed to the characteristics of the tasks; Ad Performance Estimation involves predicting numbers in the range , and A3 Recognition requires selecting all suitable labels from more than 20 labels, suggesting that the variety of data features and outputs could not be handled by few-shot alone. Both Ad Acceptability and Ad Consistency are binary classification tasks, yet the performance gap between fine-tuned models and LLMs is more noticeable in the Ad Acceptability task, with a difference as large as 0.2 points, compared to the relatively smaller difference observed in the Ad Consistency task. The task with the smallest difference was Ad Similarity, with a difference of only 0.06 to 0.10.
@@
PLMs vs. Human. Human evaluators outperform models in Ad Acceptability and Ad Consistency. In both tasks, models tends to have high accuracy and a low F1-score, while humans exhibit the opposite trend. Particularly in the Ad Consistency task, where the label distribution is unbalanced, even the best model achieves an F1-score of only 0.55, whereas human performance reaches 0.8. This suggests that humans can make better predictions in both precision and recall in an unbalanced data. On the other hand, fine-tuned models outperform humans on the A3 Recognition and Ad Similarity tasks. In the A3 Recognition task, human evaluators struggled with the diversity of output labels, similar to LLMs. However, the difference between the F1-micro and F1-macro scores is relatively small at 0.03 points for the human result. In contrast, for the fine-tuned PLMs, the difference ranges from 0.08 to 0.20 points. This indicates that human evaluators have a strong ability to generalize and can maintain higher performance even when a label appears infrequently. In the Ad Similarity task, in addition to fine-tuned models, GPT-4 outperforms human evaluators. This is the only task where LLM outperforms humans, suggesting that GPT-4 has a high level of alignment with humans in terms of semantic similarity and numerical understanding.
@@
Error Analysis on Ad Acceptability and Ad Consistency task. We have conducted a detailed analysis of two tasks, Ad Acceptability and Ad Consistency, where the model has not yet outperformed human evaluators. We analyzed three types of errors: human incorrect and model correct (IC-C), human correct and model incorrect (C-IC), and both incorrect (IC-IC), as shown in Table 7. In both tasks, the model more often predicted False labels than True labels for both IC-C and C-IC errors.. In contrast, humans provided more True labels in both correct and incorrect cases. This suggests that humans exhibit a relatively higher degree of leniency compared to models, which tend to be overly cautious when making decisions in Ad Acceptability and Ad Consistency tasks.
@@
Case Study on the Model’s Behavior on Ad Similarity Task. We used integrated gradients [18] to visualize the attention of Tohoku BERT, aiming for a deeper understanding of the model’s behavior, which can potentially enhance its performance. Figure 5 shows examples of low and high errors between predictions and the ground truth in the Ad Similarity task. In the first example, the model fails to pay sufficient attention to “online seminar,” which has a similar meaning to “webinar.” However, high attention to the date expression “9/9 (Thursday) at 13:00” allows the model to predict a score close to the correct answer. However, in the second example, the model did not pay enough attention to “English conversation coaching,” which has a similar meaning to “gaining English skills,” resulting in an incorrect prediction of dissimilarity. As we found many such cases apart from the ones mentioned, we believe that the model tends to prioritize surface-level information, especially for entities such as dates, times, and numbers. This may prevent the model from correctly identifying cases that are semantically similar even when the surface-level information differs, and vice versa.
@@
Another instance where we observed shortcomings in the models is in reasoning. For example, the models mistakenly interpret that phrases such as “half price” and “buy 2 get 1 free” have the same meaning, whereas humans easily comprehend. Additionally, pairs of phrases like “Available 24 hours a day, 7 days a week” and “Closed every Wednesday,” and “Property located near the station” and “20-minute walk from the station” often confuse the models in understanding. These types of paraphrasing are commonly used in the ad creation process; however, it is challenging to automatically capture them in a natural dataset. Therefore, a specialized dataset focusing on these phenomena is necessary to enable models to capture them accurately.
@@ 6 Related Work
Mita et al. [10] proposed an ad text generation benchmark that aims to evaluate NLG models (Step 2 in Figure 1) with surface- and semantic-level overlap metrics, including BLEU [19], ROUGE [20], and BERTScore [21], along with simple heuristics, including keyword insertion rate [22] and text length. However, this approach is inadequate for real-world applications because: (1) it requires reference texts to evaluate generated texts, making it difficult to evaluate new ads that have no references, and (2) simple heuristics cannot guarantee the quality of the ad text for delivery, as more specialized criteria are required in the actual production, such as appropriate wording, effective appeals, consistency between ad text and product information, and high predicted performance.
CTR is widely used to predict ad performance and make marketing strategies in advance. The CTR prediction task was first studied by Robinson et al. [23] and has been actively researched by Rosales et al. [24], McMahan et al. [25], Chapelle et al. [26], Yan et al. [27] and Kumar et al. [28] with traditional machine learning techniques such as logistic regression-based and factorization machines (FM)-based models [29] in past decades [14]. In recent years, neural network (NN)-based models have performed successfully, such as the CNN-based model [30] by Niu and Hou [13], LSTM-based model [31] by Gharibshah et al. [12], and the combination of FM and NN model by Guo et al. [32]. However, while there are many methods for CTR prediction, there are few studies that evaluate the quality of ad texts at a more granular level. Coarse-level feedback may be insufficient to improve ad quality, as it is unclear where edits or what changes should be made. AdTEC, on the other hand, provides more detailed evaluation results, such as acceptability and consistency, making it easier to integrate into actual AdOps workflows.
@@ 7 Conclusion
We defined five tasks to verify the quality of ad texts and built the AdTEC, a large, versatile, and comprehensive benchmark of advertising data constructed for the NLP community, based on real-world AdOps workflows. We conducted evaluations with both PLMs and human evaluators on AdTEC to explore its characteristics, offering insights into practical workflow applications and potential areas for future improvement and research. Our findings suggest that sampling directly from real data generally benefits the model, highlighting the importance for tasks centered on natural language inference and semantic understanding. We hope that the combination of our defined tasks and datasets in this paper will advance research in ad text evaluation, bridging the fields of advertising and NLP, and paving the way for new discoveries and applications.
Appendix A Dataset Documentation
@@
We adhere to the existing dataset documentation framework proposed by Gebru et al. [33] to provide comprehensive information about our dataset.
@@ A.1 Motivation
The authors of this paper created the dataset and conducted the research as part of a joint program between CyberAgent and the Nara Institute of Science and Technology for CyberAgent.
@@
This research was conducted in the joint research between CyberAgent and the Nara Institute of Science and Technology.
@@ A.5 Uses
Yes. The dataset has been used to evaluate the quality of ad texts at CyberAgent.
@@ A.7 Maintenance
Appendix B Limitations
Language
The language is limited to Japanese as the advertising agency company is located in Japan. Expressions related to online advertising vary according to language and culture. In Japan, for example, there is a cultural preference for cluttered designs [34], resulting in a lot of information being scattered around compared to the clean LPs in the US and Europe. The search queries are also different: people ask “Which car insurance is the cheapest?” in English-speaking countries, whereas short phrases, such as “cheap car insurance,” are commonly used in Japan. This indicates that different strategies and models are required for different languages and countries, and therefore, benchmarks based on different languages need to be established.
Appendix C Additional Information in Dataset Construction
C.1 Code and Data Availability
The code and data are available at https://github.com/CyberAgentAILab/AdTEC.
C.2 Annotation Guidelines
C.3 Label Distributions
@@
Appendix D Additional Information in Experiment
D.1 Model Details
The versions and URLs of the models we used are listed in Table 8. The detailed settings for the pretraining of fine-tuned models, such as the tokenizer and dataset, are provided in Table 9.
| Name | URL |
| Tohoku BERT | tohoku-nlp/bert-{base,large}-japanese-v2 |
| XLM-RoBERTa | xlm-roberta-{base,large} |
| CALM2 | cyberagent/calm2-7b-chat |
| ELYZA | elyza/ELYZA-japanese-CodeLlama-7b-instruct |
| OpenAI | 2023-08-01-preview |
| XLM-RoBERTaBASE | — | Unigram LM | Multilingual CC | 250K | 278M |
| XLM-RoBERTaLARGE | — | Unigram LM | Multilingual CC | 250K | 559M |
D.2 Hyperparameters
We used PyTorch [35] and HuggingFace [36] libraries for the model implementation. We present the hyperparameters employed during the training of both fine-tuned encoder models and zero-/few-shot settings on LLMs. Table 12 details the specific hyperparameters used in our experiments.
| Parameter | Value |
| Fine-tuned Model | |
| Lerning Rate | {2e-5, 5.5e-5, 2e-6} |
| seed | 0 |
| Epoch | 30 |
| Early Stopping Patience | 3 |
| Optimizer | Adam |
| Max Sequence Length | 128 |
| Large Language Model | |
| Attempts Per an Instance | 5 |
| Temperature | 0.8 |
| Max New Tokens | 64 |
D.3 Prompts
The template for the prompt in each task is illustrated in Figure 7. We begin by outlining the purpose and overview of the task, followed by an explanation of the criteria and standards for evaluation. Next, we provide instructions on the response format, and in some cases, additional few-shot examples are included. Tables 14 and 15 show the prompts actually used in the tasks. Terms in curly brackets are variables to be filled in. All variables, except {few_shot_examples}, correspond to the fields described in Section 3. Note that all prompts are translated from Japanese into English for the presentation brevity.

### v2

@@ 5 Results and Discussion
Table 6 provides an overview of the results. For the complete result including base size encoder models and GPT-3.5, please refer to Appendix Table 12.
Fine-tuned Setting with Encoder Models.
@@
XLM-RoBERTa and Tohoku BERT achieved the highest or competitive scores in two or more tasks. In addition, the LARGE model outperformed the BASE model in most tasks, suggesting that an increased parameter size plays a key role in understanding ad expressions.
Zero-/Few-shot Setting with LLMs.
@@
GPT-4 achieved high performance across all three tasks, while ELYZA7b performed the best among the LLMs in Ad Consistency. A substantial difference was observed between the open LLMs and OpenAI’s LLMs in the Ad Similarity task. CALM27b and ELYZA7b scored close to 0, indicating no correlation, whereas OpenAI’s models, especially GPT-4, achieved competitive scores of 0.776/0.811. Thus, the open LLMs used in this study struggle to handle semantic similarity or numerical answers, whereas GPT-4 performs considerably better. In the Ad Performance Estimation task, all LLMs produced uncorrelated responses close to 0, indicating that accurately predicting ad performance remains a challenge for open LLMs.
Fine-tuned Encoder Models vs. Zero-/Few-shot LLMs.
@@
Overall, the fine-tuned models outperformed the LLMs. The difference was substantial, ranging from 0.2 to 0.6, particularly for the Ad Acceptability, Ad Performance Estimation, and A3 Recognition tasks. This can be attributed to the characteristics of the tasks; Ad Performance Estimation involves predicting numbers in the range , and A3 Recognition requires selecting all suitable labels from more than 20 labels, suggesting that the variety of data features and outputs could not be handled by few-shot alone. Both Ad Acceptability and Ad Consistency are binary classification tasks, yet the performance gap between fine-tuned models and LLMs is more noticeable in the Ad Acceptability task, with a difference as large as 0.2 points, compared to the relatively smaller difference observed in the Ad Consistency task. The task with the smallest difference was Ad Similarity, with a difference of only 0.06 to 0.10.
PLMs vs. Humans.
@@
Human evaluators outper-form models in Ad Acceptability and Ad Consistency. In both tasks, models tends to have high accuracy and low F1-score, while humans exhibit the opposite trend. The evaluations by PLMs demonstrate high precision but low recall in both tasks, resulting in a low F1-score. According to Table 11, both datasets exhibit an imbalanced label distribution. Models tend to predict the majority label, which likely leads to high accuracy. In contrast, human experts achieve a balanced precision and recall, demonstrating robustness against the imbalanced label distribution and suggesting a higher level of generalization performance compared to the PLMs. Particularly in the Ad Consistency task, where the label distribution is unbalanced, even the best model achieves an F1-score of only 0.55, whereas human performance reaches 0.8. This suggests that humans can make better predictions in both precision and recall on unbalanced data. Meanwhile, fine-tuned models outperform humans on the A3 Recognition and Ad Similarity tasks. In the A3 Recognition task, human evaluators struggled with the diversity of output labels, similar to LLMs. However, the difference between the F1-micro and F1-macro scores is relatively small at 0.03 points for humans. In contrast, for the fine-tuned PLMs, the difference ranges from 0.08 to 0.20 points. This indicates that human evaluators have a strong ability to generalize and can maintain higher performance, even when a label appears infrequently. In the Ad Similarity task, GPT-4, in addition to fine-tuned models, outperforms human evaluators. This is the only task where LLM outperforms humans, suggesting that GPT-4 has a high level of alignment with expert humans in terms of semantic similarity and numerical understanding.
@@
We conducted a detailed analysis of two tasks: Ad Acceptability and Ad Consistency, where the model has not yet outperformed human evaluators. We analyzed three types of errors: human incorrect and model correct (IC-C), human correct and model incorrect (C-IC), and both incorrect (IC-IC), as shown in Table 7. In both tasks, the model more frequently predicted False labels than True labels for both IC-C and C-IC errors. In contrast, humans provided more True labels in both correct and incorrect cases. This suggests that humans exhibit a relatively higher degree of leniency compared to models, which tend to be overly cautious when making decisions in the Ad Acceptability and Ad Consistency tasks.
Additionally, in the Ad Acceptability task, the model can easily identify elements that can be detected heuristically, such as “incorrect use of symbols” or grammatically incorrect or non-fluent expressions. However, it struggles with evaluating expressions that are fluent but semantically unnatural, such as “Engineer’s career / Engineer’s Job” (redundancy in the same appeal) or “Condominium Sales / Single-family Home” (contradictory appeals). To address these challenges, strategies such as adding training data or providing the LLM with more detailed judgment criteria could enhance accuracy.
Regarding the Ad Consistency task, the model generally performs well when expressions from the LP are included in the ad text. However, it tends to make errors when this is not the case. There are two potential causes for this scenario. One is the use of meta-expressions, such as the word “official.” In ad text, “official” is often used to reassure customers, but it is frequently omitted in LPs because there is no need to emphasize it. To address this issue, accuracy might be improved by further training on similar data or by incorporating a specific mechanism to infer whether the LP is “official.” The second cause involves referring to non-verbal information, such as images, in the ad text. Although this is beyond the scope of the current study, it might be addressed by leveraging multimodal models.
Case Study on the Model’s Behavior on Ad Similarity Task.
@@
We used integrated gradients Sundararajan et al. (2017) to visualize the attention of Tohoku BERT, aiming for a deeper understanding of the model’s behavior, which can potentially enhance its performance. Figure 5 shows examples of low and high errors between predictions and the ground truth in the Ad Similarity task. In the first example, the model fails to pay sufficient attention to “online seminar,” which has a similar meaning to “webinar.” However, high attention to the date expression “9/9 (Thursday) at 13:00” allows the model to predict a score close to the correct answer. However, in the second example, the model did not pay enough attention to “English conversation coaching,” which has a similar meaning to “gaining English skills,” resulting in an incorrect prediction of dissimilarity. As we found many such cases, we believe that the model tends to prioritize surface-level information, especially for entities such as dates, times, and numbers. This may prevent the model from correctly identifying cases that are semantically similar, even when the surface-level information differs, and vice versa.
@@
Another instance where we observed shortcomings in the models is in reasoning. For example, the models mistakenly interpret the phrases “half price” and “buy 2 get 1 free” as having the same meaning, whereas humans easily comprehend the difference. Additionally, pairs of phrases like “Available 24 hours a day, 7 days a week” and “Closed every Wednesday,” as well as “Property located near the station” and “20-minute walk from the station” often confuse the models in understanding. These types of paraphrasing are commonly used in the ad creation process; however, it is challenging to automatically capture them in a natural dataset. Therefore, a specialized dataset focusing on these phenomena is necessary to enable models to capture them accurately.
@@ 6 Related Work
Evaluating Ad Text Quality
Mita et al. (2024) proposed an ad text generation benchmark that aims to evaluate NLG models (Step 2 in Figure 2) with surface- and semantic-level overlap metrics, such as BLEU Papineni et al. (2002), ROUGE Lin (2004), and BERTScore Zhang et al. (2020), along with simple heuristics, including keyword insertion rate Mishra et al. (2020) and text length. However, this approach is inadequate for real-world applications because (1) it requires reference texts to assess generated texts, making it difficult to evaluate new ads that have no references, and (2) simple heuristics cannot guarantee the quality of the ad text for delivery, as more specialized criteria are required in actual production, such as appropriate wording, effective appeals, consistency between ad text and product information, and high predicted performance.
Ad text quality is typically represented by metrics of user behavior obtained after ad delivery as illustrated in Step 6 of Figure 2, and CTR is widely used to predict ad performance and make marketing strategies. The CTR prediction task was first studied by Robinson et al. (2007) and has been actively researched by Rosales et al. (2012), McMahan et al. (2013), Chapelle et al. (2014), Yan et al. (2014), and Kumar et al. (2015), alongside traditional machine learning techniques, such as logistic regression-based and factorization machine (FM)-based models Rendle (2010) over the past decades Yang and Zhai (2022). While there are many methods for CTR prediction, few studies have evaluated the quality of ad texts at a more granular level. Coarse-level feedback may be insufficient to improve ad quality, as it is unclear where edits should be made or what changes are necessary. Meanwhile, AdTEC provides more detailed evaluation results, such as acceptability and consistency, making it easier to integrate into actual AdOps workflows.
@@ 7 Conclusion
We defined five tasks to verify the quality of ad texts and developed the AdTEC, a large, versatile, and comprehensive benchmark of Japanese advertising data constructed for the NLP community, based on real-world AdOps workflows. We conducted evaluations with both PLMs and human evaluators on AdTEC to explore its characteristics, offering insights into practical workflow applications and identifying potential areas for future improvement and research. Our findings suggest that sampling directly from real data generally benefits the model, highlighting the importance of tasks centered on natural language inference and semantic understanding. We hope that the combination of our defined tasks and datasets will advance research in ad text evaluation, bridging the fields of advertising and NLP and paving the way for new discoveries and applications.
Limitations
Language.
The language was limited to Japanese, as the advertising agency is located in Japan. Expressions related to online advertising vary according to language and culture. For instance, in Japan, for example, there is a cultural preference for cluttered design Masuda and Nisbett (2001), which results in a lot of information being scattered around, compared to the clean LPs in the US and Europe. This means that different strategies and models may be required for different languages and countries. Although exploring the adaptability of datasets across languages is beyond the scope of this study, it represents an exciting opportunity for future research, especially because the field of NLP in advertising remains largely unexplored.
Additionally, existing research indicates that machine translation techniques could potentially mitigate these limitations Miyazaki and Shimizu (2016); Masad et al. (2023); Wang and Hershcovich (2023). Moreover, the methodology for designing these tasks and datasets is language-independent, making it applicable to other languages. Tasks such as assessing the quality of ad text (Ad Acceptability), checking the consistency of ad text with landing pages (Ad Consistency), and rephrasing ad texts (Ad Similarity) can enhance operational efficiency across different languages. By referencing the motivations and methods outlined in this paper, future researchers and engineers can design similar tasks without starting from scratch when working in different languages.
Finally, it is important to note that the primary contribution of our study is the definition of real-world advertising operations as distinct tasks and the integration of these tasks into a comprehensive benchmark, using the Japanese dataset as a case study.
Image Data.
This study focuses solely on textual data, with image data excluded from the scope for the following reasons: (1) search engine ads consist only of text data, (2) LP images often convey content through text, (3) the performance improvement gained is small compared to the effort required to handle image data. Thus, we consider image data as supplementary in this study. While certain types of ads, such as display ads, primarily use images, further exploration of this area would be a promising direction for future work.
Data Bias.
Since the clients are limited to those who have contracts with the advertising agency, the dataset may not entirely represent all industries. Consequently, the findings of this study may not be fully generalizable to the broader internet advertising market. Future research could benefit from including a wider range of data sources to address these limitations.
Ethical Consideration
Our AdTEC dataset comprises documents such as ad texts, LP texts, keywords, and industry types. The dataset includes only a subset of the original data, sampled over a specific period and refined through preprocessing, as detailed in §3.2 of our paper.
Each instance in the dataset consists of ad text data along with related information, and each instance is labeled, as discussed in the task design section (§3). While the dataset is self-contained, it does not include any confidential information, as all data intended for internal use were made public through appropriate channels. However, it is important to note that the use of automatically generated ad texts by NLG models may introduce unnatural expressions, which could be considered a source of noise.
The data were collected in 2023 using both manual human curation and software programs. Authors and other full-time employees were involved in the data collection process. The were compensated at rates above the minimum wage and provided with basic social security benefits. Although no formal ethical review was conducted, individuals involved were directly notified and their consent was obtained before data collection..
Preprocessing, cleaning, and labeling were performed on the data, and the raw data were saved for potential future uses. This dataset has already been used to evaluate ad text quality at the author’s company and could be utilized for tasks such as language modeling on acceptable ad texts and hallucination detection. The dataset will be distributed to third parties via a URL and will be available under the Creative Commons Attribution-NonCommercial-ShareAlike 4.0 (CC-BY-NC-SA 4.0) International License. The authors will support and maintain the dataset, and updates may be made to correct errors or delete instances upon request. Users will be informed of any updates through our website or codebase.
Acknowledgements
We would like to thank Yoshiko Shinozuka and other members of the KiwamiTD444https://www.cyberagent.co.jp/news/detail/id=24670 team at CyberAgent for their valuable comments and assistance with the background research and data collection.
Appendix A Datasheets
@@
We adhere to the existing dataset documentation framework proposed by Gebru et al. (2021) to provide comprehensive information about our dataset.
@@ A.1 Motivation
The authors of this paper created the dataset and conducted the research.
@@
N/A
@@ A.5 Uses
Yes. The dataset has been used to evaluate the quality of ad texts at the author’s company.
@@
| Name | URL |
| Tohoku BERT | tohoku-nlp/bert-{base,large}-japanese-v2 |
| XLM-RoBERTa | xlm-roberta-{base,large} |
| CALM2 | cyberagent/calm2-7b-chat |
| ELYZA | elyza/ELYZA-japanese-Llama-2-7b-instruct |
| OpenAI | 2023-08-01-preview |
@@ A.7 Maintenance
| XLM-RoBERTaBASE | — | Unigram LM | Multilingual CC | 250K | 278M |
| XLM-RoBERTaLARGE | — | Unigram LM | Multilingual CC | 250K | 559M |
| Parameter | Value |
| Fine-tuned Encoder Models | |
| Learning Rate | {2e-5, 5.5e-5, 2e-6} |
| Seed | 0 |
| Epoch | 30 |
| Early Stopping Patience | 3 |
| Optimizer | Adam |
| Max Sequence Length | 128 |
| Zero-/Few-shot LLMs | |
| Attempts Per an Instance | 5 |
| Temperature | 0.8 |
| Max New Tokens | 64 |
| Fine-tuned LLMs | |
| LoRA alpha | 16 |
| LoRA dropout | 0.1 |
| Bottleneck r | 16 |
Appendix B Additional Information in Dataset Construction
B.1 Annotation Guidelines
B.2 Label Distributions
@@
Appendix C Additional Information in Experiment
C.1 Model Details
C.2 Hyperparameters
We used PyTorch Paszke et al. (2019) and HuggingFace Wolf et al. (2020) libraries for the model implementation. We present the hyperparameters employed during the training of both fine-tuned encoder models and zero-/few-shot settings on LLMs are detailed in Table 10.
| Fine-tuned Encoder Models | |||||
| Zero-/Few-shot LLMs | |||||
| Fine-tuned LLMs | |||||
| ELYZA7b | 0.638/0.638 | 0.692/0.694 | 0.240/0.235 | 0.379/0.280 | 0.684/0.740 |
| Human | 0.732/0.790 | 0.703/0.807 | — | 0.564/0.538 | 0.699/0.765 |
C.3 Prompts
The template for the prompt in each task is illustrated in Figure 6. We begin by outlining the purpose and overview of the task, followed by an explanation of the criteria and standards for evaluation. Next, we provide instructions on the response format, and in some cases, additional few-shot examples are included. Tables 13 and 14 show the prompts actually used in the tasks. Terms in curly brackets are variables to be filled in. All variables, except {few_shot_examples}, correspond to the fields described in Section 3. Note that all prompts are translated from Japanese into English for presentation brevity.